Resistors R1 = 4on and R2 = 60Ω are connected series and the combination is parallel Ra 36n, the combined resistors is connected to a battery of unknown voltage. If the voltage across R2 is 10V, A. What is the voltage of the battery? B. what is the power dissipated by R
Answer #1

a)

Current flowing through R1 and R2

I12=V2/R2 =10/60 =0.16667 A

Voltage across R1 is

V1=I12R1 =0.16667*40 =6.6667 V

Voltage of the battery

V=10+6.6667 =16.6667 Volts

b)

Voltage across R3 is

V3=V =16.6667 Volts

Current through R3 is

I3=V/R3=16.6667/36=0.463 A

Power dissipated in R3 is

P3=I32R3=0.4632*36=7.716 Watts